Find the cube root of each of the following numbers by prime factorisation method.

 (vii) 110592

 

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student


(vii) 110592

Solution:

110592 = 2×2×2×2×2×2×2×2×2×2×2×2×3×3×3

By grouping the factors in triplets of equal factors,

110592 = (2×2×2)×(2×2×2)×(2×2×2)×(2×2×2)×(3×3×3)

Here, 110592 can be grouped into triplets of equal factors,

∴ 110592 = (2×2×2×2 × 3) = 48

Hence, 48 is cube root of 110592.




Thanks